Grade 12Algebra

Multiply and express as a mixed fraction:
5 × 6 ¾ …....

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

First convert the given mixed fraction into improper fraction.
= 6 whole 3⁄4
= 27/4
Now,
= 5 × (27/4)
= 135/4
= 33 whole 3⁄4

Thanks